OpenStudy (anonymous):

A certain shade of green paint requires 4 parts blue to 5 parts of yellow if you mix 16 quarts of blue paint with 25 quarts of yellow paint will you get the desired shade o green ? explain.

OpenStudy (anonymous):

You won't get the desired shade of green because you need to keep the ratio of blue to yellow paint the same, that is 4:5. If you want to mix 16 quarts of blue, you'd need 20 quarts of yellow since 16/20=4/5, which is the correct ratio. Similarly if you wanted 25 quarts of yellow, you'd need 20 quarts of blue, since 20/25=4/5.